# Joint Self-Supervised Video Alignment and Action Segmentation

### v1

@@ 1 Introduction
Though the past decade has witnessed remarkable progress in human activity understanding in videos, the majority of the research efforts have been invested in action recognition [8, 68, 75, 20], which categorizes simple actions in short videos. In this paper, we study the two less-explored problems, i.e., temporal video alignment (frame-to-frame assignment) [45, 16, 69], which establishes framewise correspondences between long videos recording a complex activity, and temporal action segmentation (frame-to-action assignment) [30, 79, 27], which assigns frames of long videos capturing a multi-phase activity to phase/action labels. Since acquiring per-frame annotations for supervised training is generally difficult and costly, we are interested in self-supervised approaches for video alignment and action segmentation.
@@
Motivated by the above observations, we first propose VAOT (see Fig. 1(a)), a novel self-supervised video alignment approach based on a fused Gromov-Wasserstein optimal transport formulation with a structural prior, which tackles order variations, background/redundant frames, and repeated actions in a single global alignment framework. Our single-task model trains efficiently on GPUs and requires few iterations to derive the optimal transport solution, while outperforming previous methods, including VAVA [45], on video alignment datasets. Moreover, we develop VASOT (see Fig. 1(b)), a joint self-supervised video alignment and action segmentation approach by exploring the relationship between self-supervised video alignment and action segmentation via a unified optimal transport framework. Our multi-task model performs on par with prior works on video alignment benchmarks yet establishes the new state of the art on action segmentation benchmarks. In addition, our joint model requires training and storing a single model and saves both time and memory consumption as compared to two separate single-task models. Lastly, we observe in Sec. 4.3 that, in a multi-task learning setting, action segmentation provides little boost to video alignment results, whereas video alignment increases action segmentation performance significantly. To the best of our knowledge, our work is the first to analyze the relationship between video alignment and action segmentation.
@@ 3.2.2 Self-Supervised Multi-Task Learning
where $w_{align}\geq 0$ and $w_{seg}\geq 0$ denote the weights for the video alignment loss $\mathcal{L}_{xy}$ and the action segmentation losses $\mathcal{L}_{xa}$ and $\mathcal{L}_{ya}$ respectively, $\mathcal{L}_{xy}$ is the cross-entropy loss between normalized similarities $\mathbf{P}^{xy}$ (computed between $\mathbf{X}$ and $\mathbf{Y}$) and pseudo-labels $\mathbf{T}^{xy\star}$ (obtained by VAOT), while $\mathcal{L}_{xa}$ is the cross-entropy loss between normalized similarities $\mathbf{P}^{xa}$ (computed based on $\mathbf{X}$ and $\mathbf{A}$) and pseudo-labels $\mathbf{T}^{xa\star}$ (derived by ASOT [79]) and $\mathcal{L}_{ya}$ is the cross-entropy loss between normalized similarities $\mathbf{P}^{ya}$ (computed between $\mathbf{Y}$ and $\mathbf{A}$) and pseudo-labels $\mathbf{T}^{ya\star}$ (obtained by ASOT [79]). This is in contrast with VAOT or ASOT [79], where ${\bm{\theta}}$ is solely learned by using either $\mathcal{L}_{xy}$ or $\mathcal{L}_{xa}$ and $\mathcal{L}_{ya}$ respectively. We do not back-propagate gradients through $\mathbf{T}^{xy\star}$, $\mathbf{T}^{xa\star}$, and $\mathbf{T}^{ya\star}$. We find balancing $w_{align}=w_{seg}=1$ yields good results for both video alignment and action segmentation, as seen in Sec. 4.2. Our joint model requires training and storing a single model, saving both time and memory usage as compared to two single-task models. As we observe in Sec. 4.3, in a multi-task learning setting, action segmentation offers little benefit to video alignment performance, while video alignment boosts action segmentation results substantially, establishing the new state of the art on action segmentation benchmarks.
@@ 4 Experiments
Datasets. We benchmark our VAOT and VASOT approaches for video alignment using three datasets, including monotonic datasets, i.e., Pouring [57] and Penn Action [81], and in-the-wild dataset, i.e., IKEA ASM [3]. Pouring includes videos of humans pouring liquids and Penn Action comprises of videos of humans playing sports, while IKEA ASM videos capture humans assembling furniture. All methods have the same training and validation splits. Moreover, for action segmentation evaluation, we use four datasets, including in-the-wild datasets, i.e., Breakfast [33], 50 Salads [61], and YouTube Instructions [2], and monotonic dataset, i.e., Desktop Assembly [35]. Breakfast and 50 Salads videos show cooking activities, while YouTube Instructions consists of instructional videos and Desktop Assembly includes videos of an assembly activity. All methods are trained and tested on the same set of videos. For 50 Salads, we evaluate at two action granularity levels, i.e., Mid with 19 actions and Eval with 12 actions. Finally, for datasets with many activities, i.e., Penn Action, Breakfast, and YouTube Instructions, we train and test the methods per activity and report the average results. Please see recent works [18, 25, 45, 16, 34, 35, 79] for more details.
@@
Competing Methods. We compare our VAOT and VASOT approaches against prior self-supervised video alignment methods, namely SAL [48], TCN [57], TCC [18], LAV [25], VAVA [45], and GTCC [16]. Among them, VAVA [45], which integrates an optimality prior into a classical Kantorovich optimal transport, is the closest to our VAOT approach. Also, we test our VASOT approach against previous self-supervised action segmentation methods, namely CTE [34], VTE [72], UDE [64], ASAL [43], TOT [35], UFSA [70], and ASOT [79]. Here, ASOT [79] is the single-task baseline for action segmentation, which we adopt in Sec. 3.2 for our multi-task VASOT approach.
Evaluation Metrics. To evaluate our VAOT and VASOT approaches for video alignment, we compute four metrics on the validation set, i.e., phase classification (Acc@{0.1,0.5,1.0}), phase progression (Progress), video alignment ($\tau$), and fine-grained frame retrieval (AP@{5,10,15}). Prior to that, we train the model on the training set and freeze it, and then train an SVM classifier or linear regressor on top of frozen features. Note that as mentioned in [18, 25, 45, 16], Progress and $\tau$ are only defined for monotonic datasets and hence are not computed for IKEA ASM. Also, to test our VASOT approach for action segmentation, we calculate three metrics, i.e., mean over frames (MoF), F1 score (F1), and mean intersection over union (mIoU). Before that, we train the model and obtain predicted action segments, and then perform Hungarian matching between predicted and ground truth action clusters. For detailed evaluation protocols and metrics, please refer to recent works [18, 25, 45, 16, 34, 35, 79].
@@ 4.2 Sensitivity Analysis Results
(f) Alignment loss weight $w_{align}$ (here, $w_{seg}=1$).
Figure 3: Sensitivity analysis results.
Here, we conduct sensitivity analyses on hyperparameters of VAOT and VASOT. Fig. 3 shows the results. We use Pouring in Figs. 3(a-e) and Desktop Assembly in Fig. 3(f). The results of $\epsilon$ are provided in our supplementary material.
Effect of $r$ and $\alpha$. From Fig. 3(a), Acc@1.0 remains stable and Progress shows small variation across all studied values of $r$, whereas $\tau$ is the most sensitive metric, peaking at $r=0.02$ and decreasing as $r$ increases. Similar observations can be made for $\alpha$ in Fig. 3(b), where Acc@1.0 and Progress are mostly stable, whereas $\tau$ fluctuates the most, performing the best with $\alpha=0.3$. Furthermore, we find that $r=0.02$ and $\alpha=0.3$ also work the best for the remaining datasets.
Effect of $\rho$ and $\zeta$. For $\rho$ in Fig. 3(c), it can be seen that Acc@1.0 and Progress remain mostly stable, whereas $\tau$ varies the most, performing the best at $\rho=0.35$. Similarly, for $\zeta$ in Fig. 3(d), we observe that $\tau$ is the most sensitive metric, peaking at $\zeta=0.5$, whereas Acc@1.0 and Progress remain steady across the analyzed value range of $\zeta$. Moreover, we notice that $\rho=0.35$ and $\zeta=0.5$ also yield the best results for the remaining datasets.
Effect of $w_{seg}$ and $w_{align}$. We study the relationship between action segmentation and video alignment in our multi-task VASOT approach by varying their weights $w_{seg}$ and $w_{align}$ in Eq. 7. In particular, we report alignment results with various $w_{seg}$ values (while keeping $w_{align}=1$) in Fig. 3(e), and segmentation results with various $w_{align}$ values (while keeping $w_{seg}=1$) in Fig. 3(f). It can be seen from Fig. 3(e) that alignment results drop as $w_{seg}$ increases and the best overall alignment performance is obtained with $w_{seg}=1$. In contrast, it is clear from Fig. 3(f) that segmentation results improve as $w_{align}$ increases and become steady after $w_{align}$ reaches $1$. Therefore, in a multi-task learning setup, action segmentation provides little boost to video alignment results, whereas video alignment increases action segmentation performance notably. Moreover, with balancing $w_{seg}=w_{align}=1$, VASOT achieves good results for both video alignment and action segmentation. For the next section, we set $w_{seg}=w_{align}=1$ for VASOT.
@@ 4.3 State-of-the-Art Comparison Results
Video Alignment Comparison Results. We now benchmark our VAOT and VASOT approaches against previous self-supervised video alignment methods and present quantitative results in Tab. 2. Firstly, it is evident from Tab. 2 that our VAOT approach achieves the best overall performance across all datasets, outperforming all competing methods, i.e., SAL [48], TCN [57], TCC [18], LAV [25], VAVA [45], and GTCC [16]. Especially, VAOT shows major improvements over VAVA [45] on the in-the-wild IKEA ASM dataset. The results confirm the advantage of our FGW formulation with a structural prior over the classical Kantorovich formulation with an optimality prior in VAVA [45]. Next, Fig. 4 shows some qualitative results, where VAOT retrieves all 5 correct frames with the same action (Hands at shoulder) as the query image, while LAV [25] obtains 3 incorrect frames (Hands above head), highlighted by red ovals. Lastly, similar to Fig. 3(e), we find that action segmentation offers little benefit to video alignment in a multi-task learning setup, and our multi-task VASOT approach performs mostly similarly to our single-task VAOT approach in Tab. 2. This is likely because video alignment is a more complex problem involving finer-grained frame-to-frame assignment, as compared to coarser-grained frame-to-action assignment in action segmentation. Nevertheless, VASOT obtains mostly favorable results over prior works. Deep supervision [41, 42] has the potential to improve VASOT by placing the more complex video alignment loss at a deeper layer, which remains our future work. Please see our supplementary material for multi-action evaluation.
Action Segmentation Comparison Results. We test our VASOT approach against state-of-the-art self-supervised action segmentation methods and include quantitative results in Tab. 3. Firstly, from Tab. 3, VASOT consistently achieves the best results across all metrics and datasets, outperforming all competing methods, i.e., CTE [34], VTE [72], UDE [64], ASAL [43], TOT [35], UFSA [70], and ASOT [79]. While our multi-task VASOT approach shows small gains over the single-task ASOT baseline [79] on Breakfast and YouTube Instructions, our improvements on 50 Salads and Desktop Assembly are substantial. The results validate the benefit of fusing video alignment with action segmentation and demonstrate that video alignment boosts action segmentation results notably in a multi-task learning setup. Moreover, Fig. 5 shows some qualitative results, where VASOT predicts segmentations which capture action boundaries more accurately and are more closely aligned with ground truth than ASOT [79]. Please refer to our supplementary material for per-video evaluation.
@@ 5 Conclusion
This paper presents a novel approach for joint self-supervised video alignment and action segmentation. We first develop a fused Gromov-Wasserstein optimal transport with a structural prior for self-supervised video alignment, outperforming prior works. Our single-task method trains efficiently on GPUs and needs few iterations to derive the optimal transport solution. Next, we extend our approach to a unified optimal transport framework for joint self-supervised video alignment and action segmentation, yielding similar video alignment yet better action segmentation results than prior works. Our multi-task method requires training and storing a single model and saves both time and memory usage. To our best knowledge, our work is the first to explore the relationship between video alignment and action segmentation. Our future works will explore deep supervision [41, 42] and other potential applications (e.g., joint keypoint matching and clustering [19, 56]).
@@ Appendix A Supplementary Material
In this supplementary material, we first provide the implementation details of our VAOT and VASOT approaches in Sec. A.1. Next, our ablation analysis results on the Pouring dataset are included in Sec. A.2, while our sensitivity analysis results of the entropy regularization weight $\epsilon$ are presented in Sec. A.3. Furthermore, our multi-action video alignment results and our per-video action segmentation results are included in Secs. A.4 and A.5 respectively. Finally, we present some qualitative results in Sec. A.6.

### v2

@@ 1 Introduction
Though the past decade has witnessed remarkable progress in human activity understanding in videos, the majority of the research efforts have been invested in action recognition [8, 69, 76, 21], which categorizes simple actions in short videos. In this paper, we study the two less-explored problems, i.e., temporal video alignment (frame-to-frame assignment), which establishes framewise correspondences between long videos recording a complex activity, and temporal action segmentation (frame-to-action assignment), which assigns frames of long videos capturing a multi-phase activity to phase/action labels. Since acquiring per-frame annotations for supervised training is generally difficult and costly, we are interested in self-supervised approaches for video alignment and action segmentation.
@@
Motivated by the above observations, we first propose VAOT (see Fig. 1(a)), a novel self-supervised video alignment approach based on a fused Gromov-Wasserstein optimal transport formulation with a structural prior, which tackles order variations, background/redundant frames, and repeated actions in a single global alignment framework. Our single-task model trains efficiently on GPUs and requires few iterations to derive the optimal transport solution, while outperforming previous methods, including VAVA [45], on video alignment datasets. Moreover, we develop VASOT (see Fig. 1(b)), a joint self-supervised video alignment and action segmentation approach by exploring the relationship between self-supervised video alignment and action segmentation via a unified optimal transport framework. Our multi-task model performs on par with prior works on video alignment benchmarks yet establishes the new state of the art on action segmentation benchmarks. In addition, our joint model requires training and storing a single model and saves both time and memory consumption as compared to two separate single-task models. Lastly, we observe in Sec. 4.3 that, in a multi-task learning setting, action segmentation provides little boost to video alignment results, whereas video alignment increases action segmentation performance significantly.
@@ 3.2.2 Self-Supervised Multi-Task Learning
where $w_{align}\geq 0$ and $w_{seg}\geq 0$ denote the weights for the video alignment loss $\mathcal{L}_{xy}$ and the action segmentation losses $\mathcal{L}_{xa}$ and $\mathcal{L}_{ya}$ respectively, $\mathcal{L}_{xy}$ is the cross-entropy loss between normalized similarities $\mathbf{P}^{xy}$ (computed between $\mathbf{X}$ and $\mathbf{Y}$) and pseudo-labels $\mathbf{T}^{xy\star}$ (obtained by VAOT), while $\mathcal{L}_{xa}$ is the cross-entropy loss between normalized similarities $\mathbf{P}^{xa}$ (computed based on $\mathbf{X}$ and $\mathbf{A}$) and pseudo-labels $\mathbf{T}^{xa\star}$ (derived by ASOT [80]) and $\mathcal{L}_{ya}$ is the cross-entropy loss between normalized similarities $\mathbf{P}^{ya}$ (computed between $\mathbf{Y}$ and $\mathbf{A}$) and pseudo-labels $\mathbf{T}^{ya\star}$ (obtained by ASOT [80]). This is in contrast with VAOT or ASOT [80], where ${\bm{\theta}}$ is solely learned by using either $\mathcal{L}_{xy}$ or $\mathcal{L}_{xa}$ and $\mathcal{L}_{ya}$ respectively. We find balancing $w_{align}=w_{seg}=1$ yields good results for both video alignment and action segmentation, as seen in Sec. 4.2. Our joint model requires training and storing a single model, saving both time and memory usage as compared to two single-task models. As we observe in Sec. 4.3, in a multi-task learning setting, action segmentation offers little benefit to video alignment performance, while video alignment boosts action segmentation results substantially.
@@ 4 Experiments
Datasets. We benchmark our VAOT and VASOT approaches for video alignment using three datasets, including monotonic datasets, i.e., Pouring [57] and Penn Action [82], and in-the-wild dataset, i.e., IKEA ASM [3]. Pouring includes videos of humans pouring liquids and Penn Action comprises of videos of humans playing sports, while IKEA ASM videos capture humans assembling furniture. All methods have the same training and validation splits. Moreover, for action segmentation evaluation, we use four datasets, including in-the-wild datasets, i.e., Breakfast [33], 50 Salads [62], and YouTube Instructions [2], and monotonic dataset, i.e., Desktop Assembly [35]. Breakfast and 50 Salads videos show cooking activities, while YouTube Instructions consists of instructional videos and Desktop Assembly includes videos of an assembly activity. All methods are trained and tested on the same set of videos. For 50 Salads, we evaluate at two action granularity levels, i.e., Mid with 19 actions and Eval with 12 actions. Finally, for datasets with many activities, i.e., Penn Action, Breakfast, and YouTube Instructions, we train and test the methods per activity and report the average results.
@@
Competing Methods. We compare our VAOT and VASOT approaches against prior self-supervised video alignment methods, namely SAL [48], TCN [57], TCC [19], LAV [26], VAVA [45], and GTCC [17]. VAVA [45], which integrates an optimality prior into a classical Kantorovich optimal transport, is the closest to our VAOT approach. Also, we test our VASOT approach against previous self-supervised action segmentation methods, namely CTE [34], VTE [73], UDE [65], ASAL [43], TOT [35], UFSA [71], ASOT [80], and HVQ [60]. ASOT [80] is the single-task baseline for action segmentation, which we adopt in Sec. 3.2 for our multi-task VASOT approach.
Evaluation Metrics. To evaluate our VAOT and VASOT approaches for video alignment, we compute four metrics on the validation set, i.e., phase classification (Acc@{0.1,0.5,1.0}), phase progression (Progress), video alignment ($\tau$), and fine-grained frame retrieval (AP@{5,10,15}). Prior to that, we train the model on the training set and freeze it, and then train an SVM classifier or linear regressor on top of frozen features. Note that as mentioned in [19, 26, 45, 17], Progress and $\tau$ are only defined for monotonic datasets and hence are not computed for IKEA ASM. Also, to test our VASOT approach for action segmentation, we calculate three metrics, i.e., mean over frames (MoF), F1 score (F1), and mean intersection over union (mIoU). Before that, we train the model, obtain predicted action segments, and perform Hungarian matching between predicted and ground truth action clusters.
@@ 4.2 Sensitivity Analysis Results
(f) Alignment loss weight $w_{align}$ (here, $w_{seg}=1$).
Figure 3: Sensitivity analysis results. Note that (a-d) are for VAOT, while (e-f) are for VASOT.
@@
Here, we conduct sensitivity analyses on hyperparameters of VAOT and VASOT. Fig. 3(d) shows the results. We use Pouring in Figs. 3(d)(a-e) and Desktop Assembly in Fig. 3(d)(f). The results of $\epsilon$ are provided in our supplementary material.
Effect of $r$ and $\alpha$. From Fig. 3(d)(a), Acc@1.0 remains stable and Progress shows small variation across all studied values of $r$, whereas $\tau$ is the most sensitive metric, peaking at $r=0.02$ and decreasing as $r$ increases. Similar observations can be made for $\alpha$ in Fig. 3(d)(b), where Acc@1.0 and Progress are mostly stable, whereas $\tau$ fluctuates the most, performing the best with $\alpha=0.3$. Furthermore, we find that $r=0.02$ and $\alpha=0.3$ also work the best for the remaining datasets.
Effect of $\rho$ and $\zeta$. For $\rho$ in Fig. 3(d)(c), it can be seen that Acc@1.0 and Progress remain mostly stable, whereas $\tau$ varies the most, performing the best at $\rho=0.35$. Similarly, for $\zeta$ in Fig. 3(d)(d), we observe that $\tau$ is the most sensitive metric, peaking at $\zeta=0.5$, whereas Acc@1.0 and Progress remain steady across the analyzed value range of $\zeta$. Moreover, we notice that $\rho=0.35$ and $\zeta=0.5$ also yield the best results for the remaining datasets.
Effect of $w_{seg}$ and $w_{align}$. We study the relationship between action segmentation and video alignment in our multi-task VASOT approach by varying their weights $w_{seg}$ and $w_{align}$ in Eq. 7. In particular, we report alignment results with various $w_{seg}$ values (while keeping $w_{align}=1$) in Fig. 3(d)(e), and segmentation results with various $w_{align}$ values (while keeping $w_{seg}=1$) in Fig. 3(d)(f). It can be seen from Fig. 3(d)(e) that alignment results drop as $w_{seg}$ increases and the best overall alignment performance is obtained with $w_{seg}=1$. In contrast, it is clear from Fig. 3(d)(f) that segmentation results improve as $w_{align}$ increases and become steady after $w_{align}$ reaches $1$. Therefore, in a multi-task learning setup, action segmentation provides little boost to video alignment results, whereas video alignment increases action segmentation performance notably. Moreover, with balancing $w_{seg}=w_{align}=1$, VASOT achieves good results for both video alignment and action segmentation. For the next section, we set $w_{seg}=w_{align}=1$ for VASOT.
@@ 4.3 State-of-the-Art Comparison Results
Video Alignment Comparison Results. We now benchmark our VAOT and VASOT approaches against previous self-supervised video alignment methods and present quantitative results in Tab. 2. Firstly, it is evident from Tab. 2 that our VAOT approach achieves the best overall performance across all datasets, outperforming all competing methods, i.e., SAL [48], TCN [57], TCC [19], LAV [26], VAVA [45], and GTCC [17]. Especially, VAOT shows major improvements over VAVA [45] on the in-the-wild IKEA ASM dataset. The results confirm the advantage of our FGW formulation with a structural prior over the classical Kantorovich formulation with an optimality prior in VAVA [45]. Next, Fig. 4 shows some qualitative results, where VAOT retrieves all 5 correct frames with the same action (Hands at shoulder) as the query image, while LAV [26] obtains 3 incorrect frames (Hands above head), highlighted by red ovals. Lastly, similar to Fig. 3(d)(e), we find that action segmentation offers little benefit to video alignment in a multi-task learning setup, and our multi-task VASOT approach performs mostly similarly to our single-task VAOT approach in Tab. 2. This is likely because video alignment is a more complex problem involving finer-grained frame-to-frame assignment, as compared to coarser-grained frame-to-action assignment in action segmentation. Nevertheless, VASOT obtains mostly favorable results over prior works.
Action Segmentation Comparison Results. We test our VASOT approach against state-of-the-art self-supervised action segmentation methods and include quantitative results in Tab. 3. Firstly, from Tab. 3, VASOT consistently achieves the best results across all metrics and datasets, outperforming all competing methods, i.e., CTE [34], VTE [73], UDE [65], ASAL [43], TOT [35], UFSA [71], ASOT [80], and HVQ [60]. While our multi-task VASOT approach shows small gains over the single-task ASOT baseline [80] on Breakfast and YouTube Instructions, our improvements on 50 Salads and Desktop Assembly are substantial. The results validate the benefit of fusing video alignment with action segmentation and demonstrate that video alignment boosts action segmentation results notably in a multi-task learning setup. Moreover, Fig. 5 shows some qualitative results, where VASOT predicts segmentations which capture action boundaries more accurately and are more closely aligned with ground truth than ASOT [80].
@@ 5 Conclusion
This paper presents a novel approach for joint self-supervised video alignment and action segmentation. We first develop a fused Gromov-Wasserstein optimal transport with a structural prior for self-supervised video alignment, outperforming prior works. Our single-task method trains efficiently on GPUs and needs few iterations to derive the optimal transport solution. Next, we extend our approach to a unified optimal transport framework for joint self-supervised video alignment and action segmentation, yielding similar video alignment yet better action segmentation results than prior works. Our multi-task method requires training and storing a single model and saves both time and memory usage. To our best knowledge, our work is the first to explore the relationship between video alignment and action segmentation. Our future works will explore deep supervision [41, 42], complex weighting in multi-task learning [29, 9], and other potential applications (e.g., joint keypoint matching and clustering [20, 56]).
@@ Appendix A Supplementary Material
In this supplementary material, we first provide the implementation details of our VAOT and VASOT approaches in Sec. A.1. Next, our ablation analysis results on the Pouring dataset are included in Sec. A.2, while our sensitivity analysis results of the entropy regularization weight $\epsilon$ are presented in Sec. A.3. Furthermore, our multi-action video alignment results and our per-video action segmentation results are included in Secs. A.4 and A.5 respectively. Finally, we present some qualitative results in Sec. A.6 and complexity comparison results in Sec. A.7.
@@ A.6 Qualitative Results
A.7 Complexity Comparison Results
We compare the complexity (in terms of model size and training time) of our multi-task VASOT approach and the separate single-task models (VAOT+ASOT) on an Nvidia 3090Ti GPU. Using a ResNet-50 encoder on Pouring, VASOT needs ($108$ MB, $116$ mins) vs. ($216$ MB, $162$ mins) of VAOT+ASOT. Using an MLP encoder on Desktop Assembly, VASOT needs ($287$ KB, $15$ mins) vs. ($571$ KB, $21$ mins) of VAOT+ASOT. The results validate that our multi-task VASOT approach saves both memory and training time as compared to the separate single-task models (VAOT+ASOT).